At some point we want to change the AUX injection on the AS table to interfere less with the normal interferometer path, and avoid 10/90 beamsplitters which produce a fair amount of ghosting. The plan is to replace the 99/1 BS whose reflection goes to AS110 and AS55, while the transmission goes to the AS camera, with a 90/10 BS as shown in the attachment. This results in ~10% less light on the PDs compared to the pre-AUX era. Between this BS and the AS camera there will be a second 90/10 BS that sends the AUX light into the IFO, so we end up with marginally less AUX power into the IFO and the same PSL power on the AS cam. We're short optics, so this has to wait until two new beamsplitters arrive from CVI.
For a series resistance of 4.5 kohm, we are suffering from the noise-gain amplified voltage noise of the Op27 (2*3.2nV/rtHz), and the Johnson noise of the two 1 kohm input and feedback resistances. As a result, the current noise is ~2.7 pA/rtHz, instead of the 1.9 pA/rtHz we expect from just the Johnson noise of the series resistance. For the present EX coil driver configuration of 2.25 kohm, the Op27 voltage noise is actually the dominant noise source. Since we are modeling small amounts (<1dB) of measurable squeezing, such factors are important I think.
[Attachment #1] --- Sketch of the fast signal path in the coil driver board, with resistors labelled as in the following LISO model plots. Note that as long as the resistance of the coil itself << the series resistance of the coil driver fast and slow paths, we can just add their individual current noise contributions, hence why I have chosen to model only this section of the overall network.
[Attachment #2] --- Noise breakdown per LISO model with top 5 noises for choice of Rseries = 2.25 kohm. The Johnson noise contributions of Rin and Rf exactly overlap, making the color of the resulting line a bit confusing, due to the unfortunate order of the matplotlib default color cycler. I don't want to make a custom plot, so I am leaving it like this.
[Attachment #3] --- Noise breakdown per LISO model with top 5 noises for choice of Rseries = 4.5 kohm. Same comments about color of trace representing Johnson noise of Rin and Rf.
Possible mitigation strategies:
I've chosen to ignore the noise contribution of the high current buffer IC that is inside the feedback loop. Actually, it may be interesting to compare the noise measurements (on the electronics bench) of the circuit as drawn in Attachment #1, without and with the high current buffer, to see if there is any difference.
This study also informs about what level of electronics noise is tolerable from the De-Whitening stage (aim for ~factor of 5 below the Rseries Johnson noise).
Finally, in doing this model, I understand that the observation the voltage noise of the coil driver apparently decreased after increasing the series resistance, as reported in my previous elog. This is due to the network formed by the fast and slow paths (during the measurement, the series resistance in the slow path makes a voltage divider to ground), and is consistent with LISO modeling. If we really want to measure the noise of the fast path alone, we will have to isolate it by removing the series resistance of the slow bias path.
Comment about LISO breakdown plots: for the OpAmp noises, the index "0" corresponds to the Voltage noise, "1" and "2" correspond to the current noise from the "+" and "-" inputs of the OpAmp respectively. In future plots, I'll re-parse these...
I will upload more details + photos + data + schematic + LISO model breakdown tomorrow to a DCC page.
How much was the osc freq of the marconi? And then how much was the resulting freq offset between PSL and AUX?
Are we supposed to see two dips with the separation of an FSR? Or four dips (you have two sidebands)?
And the distance between the dips (28MHz-ish?) seems too large to be the FSR (22MHz-ish).
[Jon, Keerthana, Sandrine]
Thu.-Fri. we continued with PRC scans using the AUX laser, but now the "scanned" parameter is the frequency of AM sidebands, rather than the frequency of the AUX carrier itself. The switch to AM (or PM) allows us to coherently measure the cavity transfer as a function of modulation frequency.
In order to make a sentinel measurement, I installed a broadband PDA255 at an unused pickoff behind the first AUX steering mirror on the AS table. The sentinel PD measures the AM actually imprinted on the light going into the IFO, making our measurement independent of the AOM response. This technique removes not only the (non-flat) AOM transfer function, but also any non-linearities from, e.g., overdriving the AOM. The below photo shows the new PD (center) on the AS table.
With the sentinel PD installed, we proceeded as follows.
The below photo shows the measured transfer function [AUX Reflection / AUX Injection]. The measurement coherence is high to ~55 MHz (the AOM bandwidth is 60 MHz). We clearly resolve two FSRs, visible as Lorentzian dips at which more AUX power couples into the cavity. The SURFs have these data and will be separately posting figures for the measurements.
With the basic system working, we attempted to produce HOMs, first by partially occluding the injected AUX beam with a razor blade, then by placing a thin two-prong fork in the beam path. We also experimented with using a razor blade on the output to partially occlude the reflection beam just before the sensor. We were able to observe an apparent secondary dip indicative of an HOM a few times, as shown below, but could not repeat this deterministically. Besides not having fine control over the occlusion of the beams, there is also large few-Hz angular noise shaking the AS beam position. I suspect from moment to moment the HOM content is varying considerably due to the movement of the AS beam relative to the occluding object. I'm now thinking about more systematic ways to approach this.
I think if the magnet fell off, we would see high DC signal, and not 0 as we do now. I suspect satellite box or PD readout board/cabling. I am looking into this, tester box is connected to ITMY sat. box for now. I will restore the suspension later in the evening.
Suspension has now been restored. With combination of multimeter, octopus cable and tester box, the problem is consistent with being in the readout board in 1X5/1X6 or the cable routing the signals there from the sat. box.
We may lost the UL magnet or LED
Seems like DTT also works now. The trick seems to be to run sudo /usr/bin/diaggui instead of just diaggui. So this is indicative of some conflict between the yum installed gds and the relic gds from our shared drive. I also have to manually change the NDS settings each time, probably there's a way to set all of this up in a more smooth way but I don't know what it is. awggui still doesn't get the correct channels, not sure where I can change the settings to fix that.
DON"T RUN DIAGGUI AS ROOT
Furthermore, I believe we are losing more than 10% of the light due to this BS. The ASDC (which is derived from AS55 PD) level is down at ~110cts as the Michelson is fringing, while it used to be ~200 cts. I will update with a power measurement shortly. But I think we should move ahead with the plan to combine the beam into the IFO's AS mode as discussed at the meeting last week.
Is the 10% specified for P-Pol or for UNP? I contacted CVI about beamsplitters, since their website doesn't list a BS1-1064-90-... option on the website. They say a R=90% beamsplitter would be a custom job. The closest stock item they got is BS1-1064-95-2025-45UNP specified at R=95% for UNPolarized beams. They were kind enough to sent me the measured transmission curves for a recent lot of these, which is attached was uploaded to the wiki [Elog Police K: NO PROPRIETARY DOCUMENTS ON THE ELOG, which is public. Put it on our wiki and put the link here]. The figure is not labeled, but according to the contact Red is S-Pol and Blue is P-Pol, which means that this one actually has R=~90% for P, pretty much what we want. We'll need to buy two of these to make the swap in the setup.
Back to your original point: There's only a BS1-1064-10-2025-45UNP on the website, so unless we got these as custom items, the R for P-Pol is probably NOT actually 10%, just somewhere between 0% and 20%
4 std cataloge item fused silica BS1-1064-95-2025-45UNP
ordered today. They will arrive no later than July 13, 2018
Aim: To find a model that trains the simulated data of Gaussian beam spot moving in a vertical direction by the application of a sinusoidal signal.
All the attachments are in the zip folder.
The simulated video of beam spot motion without noise (amplitude of sinusoidal signal given = 20 pixels) is given in this link https://drive.google.com/file/d/1oCqd0Ki7wUm64QeFxmF3jRQ7gDUnuAfx/view?usp=sharing
I tried several cases:
I added random uniform noise (ranging from 0 to 25.5 i.e. 10% of the maximum pixel value 255) using opencv to 64*64 simulated images made in the last case( https://nodus.ligo.caltech.edu:8081/40m/13972), clipped the pixel values from 0 to 255 & trained using the same network as in the previous elog and it worked well. The variation in mean squared error with epochs is given in Attachment 1 & applied signal and output of the neural network (NN) (magnitude of the signal vs time) as well as the residual error is given in Attachment 2.
I simulated images 128*128 at 10 frames/sec by applying a sine wave of frequency 0.2Hz that moves the beam spot & resized it using opencv to 64*64. Then I trained 300cycles & tested with 1000 cycles with the following sequential model:
(i) Layers and number of nodes in each:
4096 (dropout = 0.1) -> 1024 (dropout = 0.1) -> 512 (dropout = 0.1) -> 256 -> 64 -> 8 -> 1
Activation : selu -> selu -> selu -> selu -> selu -> selu -> linear
(ii) loss function = mean squared error ( I used mean squared error to easily comprehend the result. Initially I had tried log(cosh) also but unfortunately I had stopped the run in between when test loss value had no improvement), optimizer = Nadam with default learning rate = 0.002
(iii) batch size = 32, no. of epochs = 400
I have attached the variation in loss function with epochs (Attachment 3). It was found that test loss value increases after ~50 epochs. To avoid overfitting, I added dropout to the layer of 256 nodes in the next model and removed the layer of 4096 nodes.
Same simulated data as case 2 trained with the following model,
1024 (dropout = 0.1) -> 512 (dropout = 0.1) -> 256 (dropout = 0.1) -> 64 -> 8 -> 1
Activation : selu -> selu -> selu -> selu -> selu -> linear
(ii) changed the learning rate from default value of 0.002 to 0.001. Rest of the hyperparameters same.
The variation in mean squared error in attachment 4 & NN output, applied signal & residual error (zoomed) in attachment 5. Here also test loss value increases after ~65 epochs but this fits better than the previous model as loss value is less.
Since in most of the examples in keras, training dataset was more than test dataset, I tried training 1000 cycles & testing with 300 cycles. The respective plots are attached as attachment 6 & 7. Here also, there is no significant improvement except that the test loss is increasing at a slower rate with epochs as compared to the last case.
Since most of the above cases were like overfitting (https://machinelearningmastery.com/diagnose-overfitting-underfitting-lstm-models/, https://github.com/keras-team/keras/issues/3755) except that test loss is less than train loss value in the beginning , I tried implementing case 4 with the initial model of 2 layers of 256 nodes each but with Nadam optimizer. Respective graphs in attachment 8, 9 & 10(zoomed). The loss value is slightly higher than the previous models as seen from the graph but test & train loss values converge after some epochs.
I have forgot to give ylabel in some of the graphs. It's the magnitude of the applied sine signal to move the beam spot. In most of the cases, the network almost correctly fits the data and test loss value is lower in the initial epochs. I think it's because of the dropout we added in the model & also we are training on the clean dataset.
MEDM, EPICS and dataviewer seem to work, but diaggui still doesn't work (it doesn't work on Rossa either, same problem as reported here, does a fix exist?). So looks like only donatella can run diaggui for now. I had to disable the systemd firewall per the instructions page in order to get EPICS to work. Also, there is no MATLAB installed on this machine yet. sshd has been enabled.
Two out of the four over-head fluorescent lights in the X end of the interferometer were flickering today.
We (Rana and I) are re-assembling the temperature controls on the seismometer to attempt PID control and then improve it using reinforcement learning.
We tried to re-assemble the connections for the heater and in-loop temperature sensor on the can that covers the seismometer.
We fixed (soldered) two of the connections from the heater circuit to the heater, but did not manage to get the PID working as one of the wires attached to the MOSFET had come off. Re-soldering the wire would be attempted tomorrow.
Equipment for undertaking all this is still left at the X-end of the interferometer and will be cleared soon.
pianosa has been upgraded to SL7. I've made a controls user account, added it to sudoers, did the network config, and mounted /cvs/cds using /etc/fstab. Other capabilities are being slowly added, but it may be a while before this workstation has all the kinks ironed out. For now, I'm going to follow the instructions on this wiki to try and get the usual LSC stuff working.
Initial tests look promising. Local damping works and I even locked the X arm using POX, although I did it in a fake way by simply inserting a x5.625 (=2.25 kohm / 400 ohm) gain in the coil driver filter banks. I will now tune the individual loop gains to account for the reduced actuation range.
Now I have changed the loop gains for local damping loops, Oplev loops, and POX locking loop to account for the reduced actuation range. The dither alignment servo (X arm ASS) has not been re-commissioned yet...
I took this opportunity of EX downtime to change the supply voltage for the AA unit (4-pin LEMO front panel) in 1X9 from +/-5V to +/-15V. Inside the AA board are INA134 and DRV135 ICs, which are rated to work at +/-18V. In the previous state, the inputs would saturate if driven with a 2.5Vpp sine wave from a DS345 func. gen. After the change, I was able to drive the full range of the DS345 (10Vpp), and there was no saturation seen. This AA chassis is only used for the OSEM signals and also some ALS signals. Shadow sensor levels and spectra are consistent before and after the change. The main motivation was to not saturate the Green PDH Reflection signal in the digital readout. The steps we took were:
> 2. Weighted screw rod at the bottom for tilting the mirror-holder:
Too long. The design of the holder should be check with the entire assembly.
We should be able to make it compact if we heavier weights.
How are these weights fixed on the shaft?
Also can we have options for smaller weights for the case we don't need such a range?
Note the mass of the weights.
> 3. Set-screws on both side of wire clamp to adjust its horizontal position:
How much is the range of the clamp motion limited by the slot for the side screws and the slot for the protrusion? Are they matched?
Can you show us the design of the slot made on the mirror holder?
Where is the center of mass (CoM) for the entire mirror holder assy and how much is the height gap between the CoM and the wire release points. Can you do this with 3/8" and 1/2" fused silica mirrors?
Here’s a quick summary of the Tip-Tilt Design updates (all files are in the dropbox in [TipTiltSus>TT_New]) that I have been working on with Koji and Steve's help.
1. Plate on top to hold mirror in place:
The plate is 0.5 mm thick. I did a rough FEA with 10 N force on the point of pressure on it, and it bent easily.
2. Weighted screw rod at the bottom for tilting the mirror-holder:
I did a very simplified free body analysis to calculate the required length of the rod to achieve a +/- 15 mRad tilt, and got around 1.5 inches.
3. Set-screws on both side of wire clamp to adjust its horizontal position:
1. Used the same screw size in most places to reduce complexity.
2. The mirror holder I have worked on is a little different from the actual piece I have on my table. Which one do you prefer (Koji)?
(Jon, Keerthana, Sandrine)
We tried to align the AUX and PSL laser yesterday. We collected the data from the spectrum analyser for the Y-ARM reflection and also for the Y-ARM transmission from the ETM mirror. I am attaching the plots here.
[Jon, Keerthana, Sandrina]
Yesterday we carried out preliminary proof-of-concept measurements using the new AS-port-injected AUX laser to resolve cavity mode resonances.
At the time we started, I found the beat note levels consistent with what Johannes had reported the night before post-realignment. Hence we did not change the AUX alignment.
Test 1: YARM Mode Scan
Test 2: PRC Mode Scan
The SURFs have the data from last night's scans and will be separately posting plots of these measurements. We'll continue today with mode scans using AM sidebands rather than the AUX RF offset.
I decided to take a quick look at the data. Changes made to the ETMX coil driver board:
I also took the chance to check the integrity of the LM6321 ICs. In the past, a large DC offset on the output pin of these has been indicative of a faulty IC. But I checked all the ICs with a DMM, and saw no anomalies.
Measurement condition was that (i) the Fast input was terminated to ground via 50ohm, (ii) the Bias input was shorted to ground. SR785 was used with G=100 Busby preamp (in which Steve installed new batteries today, for someone had left it on for who knows how long) for making the measurement. The voltage measurement was made at the D-Sub connector on the front panel which would be connected to the Sat. Box, with the coil driver not connected to anything downstream.
Summary of results:
[Attachment #1] - Noise measurement out to 800 Hz. The noise only seems to agree with the LISO model above 300 Hz. Not sure if the low-frequency excess is real or a measurement artefact. Tomorrow, I plan to make an LPF pomona box to filter out the HF pickup and see if the low-frequency characteristics change at all. Need to think about what this corner freq. needs to be. In any case, such a device is probably required to do measurements inside the VEA.
[Attachment #2] - Noise measurement for full SR785 span. The 19.5 kHz harmonics are visible. I have a theory about the origin of these, need to do a couple of more tests to confirm and will make a separate log.
[Attachment #3] - zip of LISO file used for modeling coil driver. I don't have the ASCII art in this, so need to double check to make sure I haven't connected some wrong nodes, but I think it's correct.
Measurements seem to be consistent with LISO model predictions.
*Note: Curves labelled "LISO model ..." are really quad sum of liso pred + busby box noise.
My main finding tonight is: With the increased series resistance (400 ohm ---> 2.25 kohm), LISO modeling tells me that even though the series resistance (Johnson noise) used to dominate the voltage noise at the output to the OSEM, the voltage noise of the LT1125 in the bias path now dominates. Since we are planning to re-design the entire bias path anyways, I am not too worried about this for the moment.
gautam noon 21 June 2018: I was looking at the wrong LISO breakdown curves. So the input stage Op27 voltage noise used to dominate. Now the Bias path LT1125 voltage noise dominates. None of the conclusions are affected... I've uploaded the corrected plots and LISO file here now.
I finished the re-soldering work today, and have measured the coil driver noise pre-Mods and post-Mods. Analysis tomorrow. I am holding off on re-installing the board tonight as it is likely we will have to tune all the loops to make them work with the reduced range. So ETMX will remain de-commissioned until tomorrow.
Aim: To measure the optical power from led using a powermeter.
Yesterday Gautam drilled a larger hole of diameter 5mm in the box as an aperture for led (aperture angle is approximately 2*tan-1(2.5/7) = 39 deg). I repeated the measurements that I had done before (https://nodus.ligo.caltech.edu:8081/40m/13951). The measurents of optical power measured using a powermeter and the corresponding input voltages are listed below.
So we are able to receive optical power close to the value (1.6mW) given in Thorlabs datasheet for LED1050E (https://www.thorlabs.com/drawings/e6da1d5608eefd5c-035CFFE5-C317-209E-7686CA23F717638B/LED1050E-SpecSheet.pdf). I hope we can proceed to BRDF measurements for CCD calibration.
Steve: did you center the LED ?
You should wipe off the table cover before you take it off next time.
It is important to turn up the PSL encloure HEPA Variac voltage if you are working in there. It takes less than 10 minutes to reach lab condition.
Lab air count normal. It is not logged. I have a notebook of particle count on the SP table next to the Met One counter.
Chris replaced some air condition filters and ordered some replacement filter today.
We did swap a lens as discussed in elog 13968, but they both had f=100mm specified, the difference being one was AR-coated for 1064 and bi-convex, while the other one was plano-convex and had a different coating. The reason for the large beam spot was something else: The fiber wasn't sitting in the coupler properly. When reconnecting the fiber after taking it out make sure to align the key on the fiber end with the notch in the coupler before tightening. After discovering this the following was done:
Before leaving I restored the XARM alignment. SRM remains misaligned, LSC off. Alignment shouldn't change drastically over night, so I suggest when picking this work up tomorrow to directly look for the beats after phaselocking AUX and PSL
Per discussion today eve, barring objections, I will do the following tomorrow morning:
Not much progress today with the AUX cavity scans. I've determined there still are some alignment issues.
At the start of today a large AUX/PSL beat note was visible on the AS110 sensor, at a similar power as where we left off last night (-60 dBm). Proceeding from there, I attempted to reproduce Johannes' measurement of the cavity transmission resonances. I misaligned the X-arm, locked the Y-arm cavity, and scanned the AUX RF offset approximately 8 MHz in 2 kHz steps. This should have swept through two FSRs, but nothing was visible.
Further inspection revealed that none of the PSL light was making it back to through the AUX fiber to the PSL table. I take this to mean that the beam seen earlier on AS110 was the ITMY reflection, and that the AUX injection axis was no longer reaching ETMY. I also found that the AUX beam size just after the 90/10 beasmsplitter looks anomolously large. Maybe a lens was recently changed? In any case, the mode-matching looks like it is going to need to be readjusted.
POP QPD checkout:
The actuator (pendulum) gains for the Beam Splitter and the two ITMs were measured to be:
BS: 9.54 +/- 0.05 nm/ct [100 ohm series resistor in coil driver board]
ITMX: 2.44 +/- 0.01 nm/ct [400 ohm series resistor in coil driver board]
ITMY: 2.44 +/- 0.02 nm/ct [400 ohm series resistor in coil driver board]
Counts here refers to DAC counts at the output of the coil filter banks (as opposed to counts at the LSC servo output). The dominant (systematic) uncertainty (which isn't quoted here) in this measurement is the determination of the peak-to-peak swing of the dark port sensor, AS55_Q. I estimate this error to be ~1ct/33cts = 3%. These values are surprisingly consistent with one another once we take into account the series resistance.
The last time this was done, we used ASDC to do the measurement. But I don't know what signal conditioning ASDC undergoes (in PD or in readout electronics). In any case, in my early trials yesterday, ASDC was behaving unpredictably. So I decided to do redo the measurement.
[Attachment #1]- Flowchart describing the calibration procedure.
[Attachment #2] - AS55_Q output while the Michelson was freeswinging. I had first aligned the ITMs using ASS. The peak-to-peak value of this corresponds to . So we know AS55_Q in terms of cts/m of MICH displacement.
[Attachment #3] - Magnitudes of transfer function from moving one of the MICH optics, to the now calibrated AS55_Q. Fits are to a shape , with being the fitted parameter. Coherence during the measurement is also plotted.
 - http://www.phys.ufl.edu/~bernard/papers/CQG20_S903.pdf
Of course, many (but no all) of the optics were custom-ordered back in ~2000.
Yesterday, I moved the following optics:
After moving these components around a bit, I locked them down once I was happy that the beam was pretty well centered on both of them, and also on AS110 and AS55 (measured using O'scope with single bounce from one ITM, other optics misaligned).
The beam was close to clipping on the lens mentioned in #1, probably because this wasn't checked when the 90-10 BS was installed for the AUX laser. Furthermore, I believe we are losing more than 10% of the light due to this BS. The ASDC (which is derived from AS55 PD) level is down at ~110cts as the Michelson is fringing, while it used to be ~200 cts. I will update with a power measurement shortly. But I think we should move ahead with the plan to combine the beam into the IFO's AS mode as discussed at the meeting last week.
Unrelated to this work, but c1psl and c1iscaux were keyed.
ASDC has something weird going on with it - my main goal yesterday was to calibrate the actuators of ITMX, ITMY and BS using the Michelson. But with the Michelson locked on a dark fringe, the ASDC level changed by up to 50 counts seemingly randomly (bright fringe was ~1000 cts, I had upped the whitening gain to +21dB), even though the CCD remained clearly dark throughout. Not sure if the problem is in the readout electronics or in the PD itself.
The initial local backup with rsync was done. Now the new 4TB disk is (supposed to be) automatically mounted at boot as /media/40mBackup so that we can run the daily backup on this disk. (<- This was confirmed by "sudomount -a")
controls@chiara|~> sudo blkid
/dev/sde1: UUID="92dc7073-bf4d-4c58-8052-63129ff5755b" TYPE="ext4"
controls@chiara|~> cat /etc/fstab
UUID=92dc7073-bf4d-4c58-8052-63129ff5755b /media/40mBackup ext4 defaults 0 0
controls@chiara|~> sudo blkid
/dev/sde1: UUID="92dc7073-bf4d-4c58-8052-63129ff5755b" TYPE="ext4"
controls@chiara|~> cat /etc/fstab
UUID=92dc7073-bf4d-4c58-8052-63129ff5755b /media/40mBackup ext4 defaults 0 0
Here I've used UUID rather than the device name "/dev/sde1" because the device name can get altered depending on the order of the usb connection.
This new disk is just a bare HDD drive sitting on the top of the chassis. We eventually want to accommodate this disk in the chassis so that we can recover the function only with the modification of /etc/fstab. We need to wait for a next chance to have chiara down. In fact, when we can isolate chiara, we want to use this disk as the main disk and install another 4TB disk as a backup.
Since I am finishing my job at the lab, I have stored all my electronics in a box (attachment 1) and placed it under the table in the control room where some other electronics are stored. The box contains the heater circuit box, two temperature sensor boards, one temperature sensor, a short power cable and +/- 15V supply cables. In the lab I left the wires for the current setup and tied them down to the wall so that they aren't in the way (attachment 2). I left the can as is and the other temperature sensor is still attached to the inside of the can. I have labeled the wires going from the sensor as 'in' and 'out'. I've also left the wires for the heater there as well (attachment 3). I turned off the PID control and deactivated the tmux session on megatron.
Thanks to Rana and the LIGO team for giving me the opportunity to work at the 40m on this project with the seismometer.
I'm running a comsol job on optimus in a tmux session named cryocavs. Should be done in less than 24 hours, judging by past durations.
I have connected a 4TB disk to chiara via a USB-SATA adapter. This disk has been recognized as /dev/sde. A GUID Partition Table (GPT), not MBR was made with gdisk to make a partition with the size beyond 2TB.
I tried to use "dd" to copy /home/cds (/dev/sdb1) to /dev/sde1, but failed. The copy was done (taking ~12h) and the partition was not recognized as a complete filesystem.
So I decided to use rsync instead.
sudo mkfs -t ext4 /dev/sde1
sudo mkdir /media/usb4g
sudo mount -t ext4 -o rw /dev/sde1 /media/usb4g
sudo rsync -a --progress /home/cds/ /media/usb4g
14:33 Copied 33G/1831G
14:38 Copied 36G/1831G
17:02 Copied 365G/1831G (~2.2GB/min)
01:18 Copied 1449G/1831G (~2.2GB/min)
> sent 1907955222607 bytes received 126124609 bytes 37010956.31 bytes/sec
> total size is 1907271994803 speedup is 1.00
Here's a Finesse modeling of what we're expecting to observe with this test. It uses Gautam's base model of the 40m IFO with appropriate modifications for the needed configuration.
The idea is to lock the IFO in the SRMI configuration, with the phase-locked AUX beam injected from the AS port. The AUX beam is imprinted with AM sidebands and slightly misaligned relative to the SRC so as to transfer power into HOM1. The RF network analyzer provides the drive signal for the AOM, and its frequency is swept to coherently measure the transfer function [reflected AUX beam / drive]. The reflected AUX beam is sensed by the AS110 PDA10CF.
It is also possible to drive PM sidebands as Koji suggests, but the squeezer group has encouraged using AM for practical advantages. The SNR with AM is a bit higher (less power lost into harmonics at large modulation index), there is a broadband AOM already available aligned to the SQZ beam at LLO, and there is also concern that driving strong PM could interfere with the SQZ control loops.
Attachment #1 shows the expected response to swept-AM in SRMI. Resolving just the FSR and the first-order mode splitting is sufficient to extract the SRC Gouy phase.
Since the 40m has not been opearted in SRMI since ~2016 (last done by Eric Q.), Gautam believes it may take some time to relock this configuration. However, the modeling indicates that we can likely obtain sufficient sensitivity in DRMI, which would allow us to proceed faster. Attachment #2 shows the expected response to swept-AM in DRMI. The PRC leakage signal turns out to be significantly smaller than the SRC reflection (a factor of ~30 in amplitude), so that the signal still retains its characteristic shape to a very good approximation. The tradeoff is a 10x reduction in SNR due to increased PSL shot noise reaching AS110.
Based on this, we should proceed with DRMI scans instead of PRMI next week.
The PRC FSR is, of course, very close to twice of our f1 moudlation frequency (11MHz x 2 = 22MHz) .
I still don't understand what response the measurement is looking for. I understood the idea of using the subcarrier as a stablized carrier to the PRC with a certain freq offset from the main carrier. I suppose what was swept was the AOM modulation frequency (i.e. modulation frequency of the AM applied to the subcarrier). If that is the case, the subcarrier seemed fixed at an arbitorary frequency (i.e. 50MHz) away from the carrier. If one of the AM sidebands hits the PRC resonance (i.e. 22, 44, 66MHz away from the main carrier), you still have the other sideband reflected back to the AS. Then the RF signal at the AS is still dominated by this reflected sideband. I feel that the phase modulation is rather suitable for this purpose.
If you are talking about ~MHz AM modulation by the AOM and scanning the PLL frequency from 1MHz to 60MHz, the story is different. And this should involve demodulation of the AS signal at the AM modulation frequency. But I still don't understand why we don't use phase modulation, which gives us the PDH type signal at the reflection (i.e. AS) port...
[Jon, Gautam, Johannes]
We did the following today:
This measurement seems like a fine candidate to trial the idea of looking for the FSRs (and in general, cavity resonances) of the PRC in the phase of the measured TFs, rather than the amplitude.
I did the measurement with the BeatMouth open today. Main changes:
So neglecting asymmetry in the branching ratio of the fiber beamsplitter, the asymmetry between the test PD optical path and the reference PD optical path is a single fiber mating sleeve in the former vs a collimator in the latter. In order to recover the expected number of 409 V/W for the Menlo PDs, we have to argue that the optical loss in the test PD path (fiber mating sleeve) are ~3x higher than in the NF1611 path (free space coupler). But at least the X and Y PDs show identical responses now. The error I made in the previously attached plot was that I was using the 20dB coupled output for the X PD measurement .
Revised conclusion: The measured optoelectronic response of the Menlo PDs at 10s of MHz, of ~130 V/W, is completely consistent with the numbers I reported in this elog. So rogue polarization is no longer the culprit for the discrepancy between expected and measured RF beatnote power, it was just that the expectation, based on Menlo PD specs, were not accurate.#2 of the linked elog seems to be the most likely, although "broken" should actually be "not matching spec".
While killing time b/w measurements, I looked on the ITMY optical table and found that the NF1611 I mentioned in this elog still exists. It is fiber coupled. Could be a better substitute as a Reference PD for this particular measurement.
I will repeat the measurement tomorrow by eliminating some un-necessary patch fiber cables, and also calibrating out the cable delays.
Aim : To develop a neural network on simulated data.
I developed a python code that generates a 64*64 image of a white Gaussian beam spot at the centre of black background. I gave a sine wave of frequency 0.2Hz that moves the spot vertically (i.e. in pitch). Then I simulated this video at 10 frames/sec for 10 seconds. Then I saved this data into an hdf5 file, reshaped it to a 1D array and gave as input to a neural network. Out of the 100 image frames, 75 were taken as training dataset and 25 as test data. I varied several hyperparameters like learning rate of the optimizer, number of layers, nodes, activation function etc. Finally, I was successful in reducing the mean squared error with the following network model:
I have attached the plot of the output of neural network (NN) as well as sine signal applied to simulate the video and their residula error in Attachment 1. The plot of variation in mean squared error (in log scale) as number of epochs increases is given in Attachment 2.
I think this network worked easily since there is no noise in the input. Gautam suggested to try the working of this network on simulated data with a noisy background.
Oplev sums of 240 days.
Since there have been various software/hardware activity going on (stack weighing, AUX laser PLL, computing timing errors etc etc), I decided to do a check on the state of the IFO.
The cabeling was cleaned up a little bit yesterday morning. The upper back side is still massy.
Using the numbers from the sensing measurement, I calibrated the measured in-loop MICH spectrum from Tuesday night into free-running displacement noise. For convenience, I used the noise-budgeting utilities to make this plot, but I omitted all the technical noise curves as the coupling has probably changed and I did not measure these. The overall noise seems ~x3 higher everywhere from the best I had last year, but this is hardly surprising as I haven't optimized anything for low noise recently. To summarize:
I will do a more thorough careful characterization and add in the technical noises in the coming days. The dominant uncertainty in the sensing matrix measurement, and hence this free-running noise spectrum, is that I haven't calibrated the actuators in a while.
I finally analyzed the sensing measurement I ran on Tuesday evening. Sensing responses for the DRMI DOFs seems consistent with what I measured in October 2017, although the relative phasing of the DoFs in the sensing PDs has changed significantly. For what it's worth, my Finesse simulation is here.
Jon spent some time trying to align the AUX beam to the SRC today, I got to the game kind of late so maybe others can add more detail.
The AUX beam that is reflected by the SRM looks terribly misshapen - it is quite elongated in vertical direction. Unfortunately I didn't snap a picture of it - anybody? It seemed at first as if this could be clipping - but after confirming the alignment of the AUX beam with the PSL output beam with aligned SRM, a slow dither of the SRM just moved the ugly pattern on the AS camera with no change to its shape - so clipping is unlikely. I'm now thinking that this is just the output beam of the fiber coupler after propagating ~15 meters to the SRM and back - even though this aspheric lens triplet coupler is supposed to be super-duper. I found that if I loosen the fiber slightly and pull it back just a bit at least the spot on the AS camera becomes nice and round - so maybe the fiber just doesn't sit well in this collimator? Not sure why that would be. I checked the fiber tip with the microscope, and while there was some gunk present, the central region and the core were clear (still cleaned using the fiber cleaning kit, which got rid of the debris). Either way, before switching to a different collimator I think we should give the Guoy phase measurement a shot - after all there was plenty of RF signal present on both AS110 and the PDA10CF placed at the YEND.
Looking for rogue beams on the AS table, I started placing some beam dumps. There was one particularly strong source of stray beams - a lens that was labeled with KPX094AR.33_F100. It became apparent after alignment efforts to the IFO had moved the AUX beam signifcantly off-center on this lens. According to the label it should have an AR coating for 1064nm, however judging by the amount of reflected light, it was certainly NOT AR-coated for 1064nm. I replaced it with a bi-convex f=100mm lens with confirmed AR-behavior.
The AUX laser is currently shuttered.
Per our Wednesday meeting, some items to work on are
All optics have been re-aligned. Jon/Johannes will elog about the work today.
Bad wording, sorry. Should have been channels in excess of ETMX controls. I'll add the others to the list as well.
Updated channel list and wiring diagram attached. Labels are 'F' for 'Front' and 'R' for - you guessed it - 'Rear', the number identifies the slot panel the breakout is attached to.
We have 6 of these boards now in cabinet E7
I wired all 32 channels going to the AA board directly to the ADC as described in the previous log. However, instead of using the old AA board and bypassing the whole circuit, I just used a breakout board as is shown in the first attachment. I put the board back in the rack and reconnected all of the cables.
The seismic BLRMs appear to be working again. A PSD of the BS seismometers is shown in attachment 2. Tomorrow I'll look at how much the ADC alone is suppressing the common mode 60 Hz noise on each of the channels.
Steve: 5 of ADC DAC In Line Test Boards [ D060124 ] ordered. They should be here within 10 days.